# and Qian Liu

### v1

@@ Abstract
Reconfigurable antennas possess the capability to dynamically adjust their fundamental operating characteristics, thereby enhancing system adaptability and performance. To fully exploit this flexibility in modern wireless communication systems, this paper considers a novel tri-hybrid beamforming architecture, which seamlessly integrates pattern-reconfigurable antennas with both analog and digital beamforming. The proposed tri-hybrid architecture operates across three layers: (i) a radiation beamformer in the electromagnetic (EM) domain for dynamic pattern alignment, (ii) an analog beamformer in the radio-frequency (RF) domain for array gain enhancement, and (iii) a digital beamformer in the baseband (BB) domain for multi-user interference mitigation. To establish a solid theoretical foundation, we first develop a comprehensive mathematical model for the tri-hybrid beamforming system and formulate the signal model for a multi-user multi-input single-output (MU-MISO) scenario. The optimization objective is to maximize the sum-rate while satisfying practical constraints. Given the challenges posed by high pilot overhead and computational complexity, we introduce an innovative tri-timescale beamforming framework, wherein the radiation beamformer is optimized over a long-timescale, the analog beamformer over a medium-timescale, and the digital beamformer over a short-timescale. This hierarchical strategy effectively balances performance and implementation feasibility. Simulation results validate the performance gains of the proposed tri-hybrid architecture and demonstrate that the tri-timescale design significantly reduces pilot overhead and computational complexity, highlighting its potential for future wireless communication systems.
Index Terms:
Reconfigurable antennas, tri-hybrid beamforming, tri-timescale, hybrid beamforming.
@@ I Introduction
Future wireless communication systems are expected to leverage ultra-wide spectrum resources in high-frequency bands, such as millimeter wave (mmWave), to accommodate the growing demand for ultra-high data rates [1]. To mitigate the severe propagation loss in these frequency bands, massive multiple-input multiple-output (mMIMO) systems or extremely large antenna arrays (ELAA) are typically employed to ensure reliable signal transmission [2]. In conventional antenna arrays, each antenna element typically exhibits a fixed and identical radiation patterns that concentrate energy towards predetermined directions, often perpendicular to the array. Within this framework, most existing studies focus on signal processing-based beamforming techniques, which adjust the phase and amplitude of signals across antenna elements to direct beam towards specific directions.
@@
However, the limitations of conventional pattern-fixed antennas are becoming increasingly apparent in emerging applications, such as ELAA and near-field communications [3]. Specifically, system performance is significantly influenced not only by array beamforming but also by the radiation gain of each individual antenna element. While antennas with narrow, highly energy-concentrated radiation patterns provide high gain within specific angular ranges, they suffer substantial performance degradation in the endfire direction due to a sharp decrease in signal strength. Conversely, antennas with broader radiation patterns provide wider coverage at the cost of reduced gain, ultimately compromising the quality of service.
@@
Reconfigurable antenna technology offers a promising solution to the challenges faced by conventional pattern-fixed antennas by enhancing adaptability [4]-[7]. Reconfigurable antennas can adapt their internal parameters to modify key operating characteristics such as radiation patterns [8]-[10], frequency responses [11], and polarizations [12], [13]. Among these, radiation pattern reconfiguration is particularly crucial for wireless communication systems, as it directly impacts both the transmitted and received signal strength. To enable flexible control over these characteristics, several hardware implementations have been developed, such as electronically steerable parasitic array radiator (ESPAR) antennas [14], frequency selective surfaces (FSS) [15], and pixel antennas [16], all of which share similar underlying principles and architectures. Typically, a reconfigurable antenna consists of an active element and multiple parasitic elements, where the connection states of parasitic elements are dynamically controlled to alter the current distribution within the antenna, thereby modifying its electromagnetic (EM) properties. This capability introduces new degrees of freedom (DoF) in the EM domain, offering the potential to overcome existing performance limitations and enable more efficient wireless communication systems.
@@
With significant advancements in reconfigurable antenna technology, integrating pattern-reconfigurable antennas into antenna arrays has gained considerable attention in recent years. For instance, the authors in [17] employ multi-port network theory and beamspace channel representation to develop an EM-based communication model for pixel antennas, which is then used to optimize antenna coding for maximizing channel gain. A similar approach is applied in [14] to construct an ESPAR antenna model, enabling analysis and optimization of ESPAR beamforming. Additionally, the authors in [18] introduce a decomposition method based on spherical harmonic functions to model radiation patterns in the EM domain. This approach decomposes the radiation pattern into a linear combination of orthogonal bases, facilitating the iterative design of radiation patterns and digital precoders. While these studies highlight the advantages of replacing conventional pattern-fixed antennas with innovative pattern-reconfigurable ones, they overlook the significant hardware costs and power consumption incurred by the increasing number of antennas and operating frequencies. To tackle these issues, existing literature commonly suggests hybrid beamforming architectures for mMIMO systems, as a power-efficient alternative to conventional fully-digital beamforming [19].
In hybrid beamforming, analog beamforming controls the beam direction in the radio frequency (RF) domain, while digital beamforming performs spatial signal processing in the baseband (BB) domain, leveraging a reduced number of RF chains and hardware-efficient phase shifters (PSs). Common hybrid beamforming architectures, including fully-connected [20], partially-connected [21], and dynamic architectures [22], strike a balance between performance and implementation cost. While recent advances in hardware and power-efficient hybrid architectures have enabled large-scale MIMO deployments for 6G and beyond, the limitations of conventional pattern-fixed antennas remain a significant challenge. To address this issue, we propose leveraging the flexibility of pattern-reconfigurable antenna technology to address the performance degradation caused by fixed radiation patterns, while utilizing hybrid beamforming to reduce power consumption without sacrificing performance.
@@
In this paper, we consider a novel tri-hybrid beamforming architecture that integrates pattern-reconfigurable antennas with both analog and digital beamforming. Specifically, this tri-hybrid architecture consists of three layers: A radiation beamformer in the EM domain to align radiation patterns with desired direction, an analog beamformer in the RF domain to enhance array gain, and a digital beamformer in the BB domain to mitigate multi-user interference. The concept of tri-hybrid architecture was first introduced in [23], [24]. While prior analyses and simulations have demonstrated the advantages of tri-hybrid beamforming in terms of spectral efficiency and energy consumption, several critical issues must be addressed before its practical implementation in future wireless systems. First, accurate acquisition of instantaneous channel state information (CSI) becomes more challenging due to the higher dimensionality of CSI and the limited number of RF chains, leading to increased pilot overhead and requiring advanced estimation algorithms. Second, optimizing the beamformers across all three domains is more complicated due to parameter coupling, discrete phase-shift constraints, and the nonlinear characteristics of reconfigurable antennas, etc.
Recently, a two-timescale framework has been proposed to address system overhead and design complexity in conventional hybrid beamforming. In this framework, the analog beamformer is optimized based on channel statistics over a long-timescale, while the digital beamformer is designed using low-dimensional instantaneous effective CSI over a short-timescale [25]-[27]. Building on this concept and leveraging the unique characteristics of the tri-hybrid architecture, we propose a novel tri-timescale framework for tri-hybrid beamforming design, effectively managing practical overhead and complexity. The main contributions of this paper are summarized as follows:
@@
  • We first develop a comprehensive mathematical model for the radiation beamformer based on multiple pattern-reconfigurable antennas, along with the corresponding virtual angular index channel in the EM domain. Building upon this model, we derive the formulation of the tri-hybrid beamforming architecture, establish the signal model for a multi-user multiple-input single-output (MU-MISO) system employing this architecture, and present the associated optimization problem.
  • For the tri-hybrid beamforming design, our objective is to maximize the sum-rate while satisfying constraints on radiation pattern energy, constant modulus and discrete phase shifts of PSs, a non-overlapping dynamic connection network, and the total transmit power budget. To effectively address the challenges of high pilot overhead and substantial computational complexity, we propose a novel tri-timescale beamforming framework, where the radiation beamformer is optimized over a long-timescale, the analog beamformer over a medium-timescale, and the digital beamformer over a short-timescale.
  • Extensive simulations are conducted to validate the capability of the tri-hybrid beamforming architecture in enhancing system performance, as well as the effectiveness of the proposed tri-timescale beamforming design algorithm in reducing pilot overhead and computational complexity.
@@ II-A Pattern-reconfigurable Antenna
Pattern-reconfigurable antenna is an advanced technology that enables dynamic adaptation of radiation patterns to enhance signal quality, mitigate interference, and improve coverage in wireless communication systems. Various hardware implementations have been explored to achieve flexible radiation pattern control, including ESPAR antennas [14], FSS [15], and pixel antennas [16]. Among these, pixel antenna technology has emerged as the most practical and versatile solution for integration into compact antenna arrays. The hardware prototypes of pixel-based reconfigurable antennas are illustrated in Fig. 1.
@@
Having demonstrated their feasibility in real-world applications, we now describe the hardware configuration and operational principle of the pixel-based pattern-reconfigurable antennas. As illustrated in Fig. 2, pixel antenna technology discretizes a continuous antenna surface into electrically small parasitic elements, called pixels, with a central radiating source driving the entire antenna. These pixels serving as fundamental building blocks, utilize low-voltage, binary-controlled pin diodes to connect or disconnect adjacent pixels. By adjusting the states of these switches, different pixel topologies are formed, which modify the current distribution across the antenna surface. This reconfiguration alters the electric and magnetic field components radiated by the antenna, enabling the generation of a diverse set of radiation patterns.
To evaluate the performance improvements brought by pattern-reconfigurable antennas, it is crucial to establish a mathematical model for both the radiation pattern and the corresponding channel characteristics. As aforementioned, pattern-reconfigurable antennas introduce an additional DoF in the EM domain, effectively expanding the information coverage dimension of conventional antennas, from a one-dimensional spatial representation to an $M$-dimensional angular information, where $M$ represents the number of uniform sampling points in the angular domain. Specifically, a single reconfigurable radiation pattern (exemplified by the red pattern in Fig. 2) is characterized over $M$ discrete directions in the EM domain and is represented by $\mathbf{f}_{\mathrm{EM}}\in\mathbb{R}^{M}$, with each element corresponding to the radiation gain in a specific direction. To model reception characteristics, we introduce the virtual angular index channel for the pattern-reconfigurable antenna, denoted as $\mathbf{h}_{\mathrm{EM}}\in\{0,1\}^{M}$, which serves as an index matrix indicating the spatial angles at which the user is located. Consequently, the effective gain in the EM domain can be expressed as $G_{\mathrm{EM}}=\mathbf{h}_{\mathrm{EM}}^{H}\mathbf{f}_{\mathrm{EM}}$, which quantifies the radiation intensity directed toward the user’s actual spatial angle. By optimizing the radiation pattern $\mathbf{f}_{\mathrm{EM}}$ to align with the angular domain channel characteristics, the correlation between them can be maximized, leading to enhanced radiation efficiency. This formulation serves as the foundation for the following theoretical analysis and optimization.
@@
Inspired by the increased design DoF introduced by pattern-reconfigurable antennas, we integrate this advanced hardware component into the conventional hybrid beamforming architecture and propose a tri-hybrid beamforming framework that incorporates beamforming across three domains. Specifically, radiation beamforming in the EM domain is employed to dynamically shape the radiation pattern, analog beamforming in the RF domain is utilized to enhance array gain, and digital beamforming in the BB domain is implemented to mitigate multi-user interference.
@@ II-B Tri-hybrid Beamforming Architecture
To provide a clear illustration of the proposed tri-hybrid beamforming architecture, we consider a mmWave MU-MISO downlink communication system, as shown in Fig. 3. Specifically, a base station (BS) equipped with a uniform linear array (ULA) serves multiple single-antenna user equipments (UEs). The BS employs reconfigurable antennas that dynamically adjust their radiation patterns, while each UE is equipped with a conventional omnidirectional antenna. The antenna spacing at the BS is set to $d={\lambda}/{2}$, where $\lambda$ is the center carrier wavelength. Let $\mathcal{N}_{\mathrm{t}}\triangleq\{1,2,\ldots,N_{\mathrm{t}}\}$ denote the set of BS antennas, $\mathcal{K}\triangleq\{1,2,\ldots,K\}$ denote the set of UEs, and $\mathcal{L}\triangleq\{1,2,\ldots,L\}$ denote the set of propagation paths between the BS and each UE. In the following, we mathematically formulate the radiation beamformer $\mathbf{F}_{\mathrm{EM}}$ in the EM domain, the analog beamformer $\mathbf{F}_{\mathrm{RF}}$ in the RF domain and the digital beamformer $\mathbf{F}_{\mathrm{BB}}$ in the BB domain, establishing the foundation for the tri-hybrid beamforming architecture.
EM domain: Building upon the radiation pattern model of a single pattern-reconfigurable antenna $\mathbf{f}_{\mathrm{EM}}$ presented in Section II-A, we extend this model to characterize the programmable radiation pattern of the entire antenna array at the BS, which consists of $N_{\mathrm{t}}$ reconfigurable antennas. The resulting radiation beamformer in the EM domain is expressed in the following matrix form
@@
where the $m$-th element of $\mathbf{f}_{\mathrm{EM},n}$ represents the radiation gain of the $n$-th antenna in the $m$-th angular direction. To ensure fairness in pattern design, the total energy of each antenna’s radiation pattern is restricted and normalized as $\|\mathbf{f}_{\mathrm{EM},n}\|^{2}=1,\forall n$. To further characterize the spatial mapping of the reconfigurable antennas, we introduce the virtual angular index channel matrix
$\overline{\mathbf{H}}_{\mathrm{EM},k}\triangleq\mathbf{I}_{N_{\mathrm{t}}}\otimes\mathbf{H}_{\mathrm{EM},k}\in\{0,1\}^{MN_{\mathrm{t}}\times LN_{\mathrm{t}}},~{}\forall k,$ (2)
@@
where $\mathbf{H}_{\mathrm{EM},k}\triangleq[\mathbf{h}_{\mathrm{EM},k,1},\mathbf{h}_{\mathrm{EM},k,2},\ldots,\mathbf{h}_{\mathrm{EM},k,L}]\in\{0,1\}^{M\times L}$ that specifies the spatial angles associated with the $k$-th UE. Each column $\mathbf{h}_{\mathrm{EM},k,l}$ contains exactly one non-zero element, indicating the angular direction of the $l$-th propagation path for the $k$-th UE. Consequently, the radiation gain $\mathbf{G}_{\mathrm{EM},k}\in\mathbb{R}^{LN_{\mathrm{t}}\times N_{\mathrm{t}}}$ of all antennas towards the $k$-th UE can be formulated as
$\mathbf{G}_{\mathrm{EM},k}$ $=\overline{\mathbf{H}}_{\mathrm{EM},k}^{H}\mathbf{F}_{\mathrm{EM}}$ (3)
$=\mathrm{blkdiag}\{\mathbf{H}_{\mathrm{EM},k}^{H}\mathbf{f}_{\mathrm{EM},1},\ldots,\mathbf{H}_{\mathrm{EM},k}^{H}\mathbf{f}_{\mathrm{EM},N_{\mathrm{t}}}\},$
@@
where the $l$-th entry of $\mathbf{H}_{\mathrm{EM},k}^{H}\mathbf{f}_{\mathrm{EM},n}$ represents the actual radiation gain of the $n$-th antenna towards the $l$-th direction for the $k$-th UE. In summary, the matrices $\mathbf{F}_{\mathrm{EM}}$ and $\overline{\mathbf{H}}_{\mathrm{EM},k}$ mathematically capture the additional degrees of freedom introduced by reconfigurable antennas, enabling a more flexible and adaptive radiation beamforming strategy in the EM domain.
@@
RF domain: We adopt a dynamic-subarray analog beamforming architecture, where each RF chain is adaptively connected to a disjoint subset of the total transmit antennas via a switch network. By fully exploiting the design flexibility of dynamic antenna selection, this architecture achieves an optimal balance between system efficiency and communication performance. Without loss of generality, we assume that the number of RF chains $N_{\mathrm{RF}}$ is not less than the number of UEs $K$, i.e., $N_{\mathrm{RF}}\geq K$, and the set of RF chains is denoted as $\mathcal{N}_{\mathrm{RF}}\triangleq\{1,2,\ldots,N_{\mathrm{RF}}\}$. The analog beamformer employs a set of low-resolution PSs, each following a constant modulus constraint with quantized phase control of $B$ bits. The discrete phase set of PSs is given by $\mathcal{F}\triangleq\{\frac{1}{\sqrt{N_{\mathrm{t}}}}e^{\jmath\frac{2\pi b}{2^{B}}}|b=0,1,\ldots,2^{B}-1\}$. To facilitate dynamic-subarray beamforming, we define the analog beamformer matrix as $\mathbf{F}_{\mathrm{RF}}\in\{\mathcal{F},0\}^{N_{\mathrm{t}}\times N_{\mathrm{RF}}}$, which implies that if the $l_{\mathrm{RF}}$-th RF chain is connected to the $n$-th antenna, the corresponding phase-shift takes a non-zero value, i.e., $\mathbf{F}_{\mathrm{RF}}(n,l_{\mathrm{RF}})\in\mathcal{F}$; if not, it is set as zero, i.e., $\mathbf{F}_{\mathrm{RF}}(n,l_{\mathrm{RF}})=0$. Moreover, to ensure non-overlapping subarrays, each antenna is dynamically assigned to only one RF chain, meaning that the analog beamformer matrix is restricted to have only one non-zero element per row, i.e., $\|\mathbf{F}_{\mathrm{RF}}(n,:)\|_{0}=1$.

### v2

@@ Abstract
Reconfigurable antennas possess the capability to dynamically adjust their fundamental operating characteristics, thereby enhancing system adaptability and performance. To fully exploit this flexibility in modern wireless communication systems, this paper considers a novel tri-hybrid beamforming architecture, which seamlessly integrates pattern-reconfigurable antennas with both analog and digital beamforming. The proposed tri-hybrid architecture operates across three layers: (i) a radiation beamformer in the electromagnetic (EM) domain for dynamic pattern alignment, (ii) an analog beamformer in the radio-frequency (RF) domain for array gain enhancement, and (iii) a digital beamformer in the baseband (BB) domain for multi-user interference mitigation. To establish a solid theoretical foundation, we first develop a comprehensive mathematical model for the tri-hybrid beamforming system and formulate the signal model for a multi-user multi-input single-output (MU-MISO) scenario. The optimization objective is to maximize the sum-rate while satisfying practical constraints. Given the challenges posed by high pilot overhead and computational complexity, we introduce an innovative tri-timescale beamforming framework, wherein the radiation beamformer is optimized over a long-timescale, the analog beamformer over a medium-timescale, and the digital beamformer over a short-timescale. This hierarchical strategy effectively balances performance and implementation feasibility. Simulation results validate the performance gains of the proposed tri-hybrid architecture and demonstrate that the tri-timescale design significantly reduces pilot overhead and computational complexity, highlighting its potential for future wireless communication systems, and revealing a practical way to exploit EM‑domain reconfigurability across three natural channel timescales.
@@ I Introduction
Future wireless communication systems are anticipated to leverage ultra-wide spectrum resources in high-frequency bands, such as millimeter-wave (mmWave), to meet the escalating demand for ultra-high data rates [1]. To counteract severe path loss in these bands, massive multiple-input multiple-output (massive MIMO) systems or extremely large antenna arrays (ELAA) have emerged as promising technologies to ensure reliable signal transmission [2]. In conventional antenna arrays, each antenna element typically possesses a fixed radiation pattern, concentrating energy toward predetermined directions, often perpendicular to the array plane. Consequently, existing research primarily focuses on signal-processing-based beamforming methods that manipulate the phase and amplitude of antenna signals across antenna elements to steer beams toward specific directions.
@@
However, the limitations of conventional antennas with fixed radiation patterns become increasingly evident in emerging applications, such as ELAA and near-field communications [3]. Specifically, system performance critically depends not only on array-level beamforming but also on individual antenna element radiation characteristics. Narrow radiation patterns with high energy concentration can deliver strong gains within specific angular sectors but experience significant performance deterioration in end-fire directions due to sharp gain reductions. Conversely, antennas with broader radiation patterns provide extended coverage at the expense of lower gain, ultimately degrading user quality of service (QoS).
@@
Reconfigurable antenna technology offers a compelling approach to addressing these limitations by introducing enhanced adaptability [4]-[9]. Such antennas dynamically adjust their internal configurations, modifying key operational characteristics, including radiation patterns [10]-[12], frequency responses [13], and polarizations [14], [15]. Among these, radiation pattern reconfiguration is particularly significant for wireless communication, directly influencing signal strength at transmission and reception. Several hardware implementations facilitating flexible control have emerged, including electronically steerable parasitic array radiator (ESPAR) antennas [16], frequency selective surfaces (FSS) [17], and pixel antennas [18]. Typically, a reconfigurable antenna comprises an active radiating element and multiple parasitic elements, whose connection states are dynamically adjusted to alter current distribution and thereby modify electromagnetic (EM) properties. This adaptability provides new degrees of freedom (DoF) in the EM domain, offering substantial potential to overcome existing performance limitations and enable more efficient wireless communication systems.
@@
Recent advances in reconfigurable antenna technology have sparked growing interest in integrating pattern-reconfigurable antennas into arrays. For instance, the authors in [19] employ multi-port network theory and beamspace channel representation to develop an EM-based communication model for pixel antennas, which is then used to optimize antenna coding for maximizing channel gain. Additionally, the authors in [20] introduce a decomposition method based on spherical harmonic functions to model radiation patterns in the EM domain. This approach decomposes the radiation pattern into a linear combination of orthogonal bases, facilitating the iterative design of radiation patterns and digital precoders. While these studies highlight the advantages of replacing conventional pattern-fixed antennas with innovative pattern-reconfigurable ones, they often neglect the growing concerns of hardware complexity and energy consumption that future wireless communication systems will encounter. Such concerns become increasingly critical in large-scale deployments, where both the antenna count and operating frequency bands continue to expand. To tackle these issues, existing literature commonly advocates hybrid beamforming architectures for mMIMO systems, as a power-efficient alternative to conventional fully-digital beamforming [21].
Hybrid beamforming architectures typically partition functionality between analog beamforming, controlling beam directions in the radio-frequency (RF) domain, and digital beamforming, performing spatial signal processing in the baseband (BB) domain. Utilizing fewer RF chains and hardware-efficient phase shifters (PSs), various hybrid architectures, such as fully-connected [22], partially-connected [23], and dynamic structures [24], achieve a balanced trade-off between performance and complexity. However, these architectures inherently rely on antennas with fixed radiation characteristics, severely restricting their adaptability to dynamic propagation environments and limiting overall system performance.
To overcome this critical limitation, leveraging pattern-reconfigurable antennas within hybrid beamforming architectures emerges as an effective solution, enabling dynamic adaptation of radiation characteristics while maintaining reduced hardware complexity and power consumption. In this paper, we consider a novel tri-hybrid beamforming architecture that integrates pattern-reconfigurable antennas with both analog and digital beamforming. This architecture can be viewed as an advanced extension of existing reconfigurable antenna-based frameworks. Specifically, the proposed tri-hybrid architecture comprises three distinct layers: an EM domain radiation beamformer aligning radiation patterns dynamically, an RF domain analog beamformer enhancing array gains, and a BB domain digital beamformer mitigating multi-user interference.
@@
The concept of this tri-hybrid architecture was first introduced in [25]-[27]. While prior analyses and simulations have demonstrated the advantages of tri-hybrid beamforming in terms of spectral efficiency and energy efficiency, several critical issues must be addressed before its practical implementation in future wireless systems. Firstly, acquiring instantaneous channel state information (CSI) becomes considerably more challenging due to increased dimensionality and limited RF chains, resulting in substantial pilot overhead. Secondly, jointly optimizing beamformers across three distinct domains introduces considerable complexity, primarily due to parameter coupling, discrete phase-shift constraints, and the nonlinear characteristics inherent to reconfigurable antennas.
To address these challenges, recent studies proposed a two-timescale beamforming framework to mitigate pilot overhead and computational complexity. Specifically, analog beamforming optimization leverages statistical CSI over a longer timescale, whereas digital beamforming employs instantaneous, low-dimensional effective CSI over a shorter timescale [28]-[30]. Inspired by this approach and capitalizing on the distinctive properties of our tri-hybrid architecture, we introduce an innovative tri-timescale beamforming framework, effectively resolving practical implementation challenges, significantly reducing channel estimation overhead, and lowering computational complexity. The main contributions of this paper are summarized as follows:
@@
  • We propose a comprehensive mathematical model for the radiation beamformer based on multiple pattern-reconfigurable antennas, including the introduction of a novel virtual angular index channel in the EM domain. Utilizing this model, we develop a tri-hybrid beamforming architecture, formulate a detailed multi-user multiple-input single-output (MU-MISO) signal model, and rigorously establish the corresponding sum-rate maximization optimization problem under realistic system constraints.
  • To overcome the inherent challenges of excessive pilot overhead and high computational complexity, we introduce an innovative tri-timescale beamforming framework. Specifically, we optimize the radiation beamformer at a long timescale according to the statistical CSI of scattering cluster cores, update the analog beamformer at a medium timescale based on high-dimensional statistical CSI of multipath rays, and adaptively design the digital beamformer at a short timescale using instantaneous low-dimensional effective CSI.
  • We conduct extensive simulations to verify the effectiveness and scalability of the proposed tri-hybrid beamforming architecture. Particularly, the proposed tri-hybrid beamforming achieves around 7dB transmit power reduction at a given sum-rate and nearly an order-of-magnitude pilot overhead reduction compared to real-time beamforming, underscoring its practical advantages for next-generation wireless systems.
@@ II-A Pattern-reconfigurable Antenna
Pattern-reconfigurable antennas represent advanced technologies that dynamically adapt their radiation patterns, significantly enhancing signal quality, mitigating interference, and improving coverage in wireless communication systems. Several hardware implementations have been explored to achieve such flexible radiation control, notably ESPAR antennas [16], FSS [17], and pixel antennas [18]. Among these, pixel antenna technology has emerged as the most practical and versatile solution, particularly suitable for integration into compact antenna arrays. Hardware prototypes of pixel-based reconfigurable antennas are illustrated in Fig. 1. This prototype employs PIN-diode switching networks that exhibit microsecond-level switching latencies and highly reliable reconfiguration behavior, confirming their suitability for rapid EM domain adaptation within the proposed framework. Moreover, these antennas typically require only a minimal number of serialized binary control lines and low-voltage direct-current (DC) bias circuits, resulting in affordable control signaling overhead and power consumption.
@@
Having established their practicality for real-world deployment, we now detail the hardware configuration and operating principles of pixel-based pattern-reconfigurable antennas. As depicted in Fig. 2, pixel antenna technology discretizes a continuous antenna surface into numerous electrically small parasitic elements, termed pixels, driven centrally by an active radiating source. These pixels, functioning as fundamental building blocks, employ low-voltage, binary-controlled PIN diodes to connect or disconnect adjacent elements. By adjusting these switches’ states, various pixel topologies are formed, modifying the current distribution across the antenna surface. This reconfiguration process effectively alters the radiated electric and magnetic field distributions, thus enabling the generation of diverse radiation patterns.
To rigorously evaluate the performance improvements brought by pattern-reconfigurable antennas, it is crucial to establish a mathematical model capturing both radiation patterns and corresponding channel characteristics. As aforementioned, pattern-reconfigurable antennas introduce an additional DoF in the EM domain, effectively expanding conventional antennas’ spatial representation from a one-dimensional model to an $M$-dimensional angular domain. As illustrated in Fig. 2, the angular space $\pi/2,\pi/2]$ is uniformly divided into $M$ sampling points, which can be expressed as $\vartheta_{m}=-\frac{\pi}{2}+\frac{\pi m}{M},~m=1,2,\ldots,M$. Specifically, each reconfigurable radiation pattern (illustrated by the red pattern in Fig. 2) is represented across $M$ discrete angular directions by a vector
$\mathbf{f}_{\mathrm{EM}}\in\mathbb{R}^{M},$ (1)
with each element corresponding to radiation gain in a particular direction. Correspondingly, we introduce the virtual angular index channel for pattern-reconfigurable antennas, denoted as
$\mathbf{h}_{\mathrm{EM}}\in\{0,1\}^{M},$ (2)
which serves as an indexing vector indicating the spatial angular information of the user. Consequently, the effective EM domain gain can be mathematically expressed as
$G_{\mathrm{EM}}=\mathbf{h}_{\mathrm{EM}}^{H}\mathbf{f}_{\mathrm{EM}}\in\mathbb{R},$ (3)
which quantifies the radiation intensity toward the user’s actual spatial angle. By optimizing the radiation pattern $\mathbf{f}_{\mathrm{EM}}$ to closely align with the angular channel $\mathbf{h}_{\mathrm{EM}}$, we can maximize their correlation, thus significantly enhancing the overall radiation efficiency. This theoretical formulation establishes a foundation for the subsequent analysis and optimization.
@@
Inspired by the increased design DoF introduced by pattern-reconfigurable antennas, we propose integrating this advanced antenna technology into conventional hybrid beamforming architectures, and develop a novel tri-hybrid beamforming framework that simultaneously considers three distinct beamforming domains. Specifically, radiation beamforming within the EM domain dynamically shapes the antenna radiation patterns, analog beamforming in the RF domain enhances array gain, and digital beamforming in the BB domain mitigates multi-user interference. This comprehensive integration fully leverages the unique advantages of each domain, thereby achieving superior performance and efficiency in complex wireless communication scenarios.
@@ II-B Tri-hybrid Beamforming Architecture
To provide a clear illustration of the proposed tri-hybrid beamforming architecture, we consider a mmWave MU-MISO downlink communication system, as shown in Fig. 3. Specifically, a base station (BS) equipped with a uniform linear array (ULA) serves $K$ single-antenna user equipments (UEs). The wireless channel between the BS and each UE comprises $L$ propagation path. The BS employs $N_{\mathrm{t}}$ reconfigurable antennas that dynamically adjust their radiation patterns, while each UE is equipped with a conventional omnidirectional antenna. The antenna spacing at the BS is set to $d={\lambda}/{2}$, where $\lambda$ is the center carrier wavelength. In the following, we mathematically formulate the radiation beamformer $\mathbf{F}_{\mathrm{EM}}$ in the EM domain, the analog beamformer $\mathbf{F}_{\mathrm{RF}}$ in the RF domain and the digital beamformer $\mathbf{F}_{\mathrm{BB}}$ in the BB domain, establishing the foundation for the tri-hybrid beamforming architecture.
EM domain: The radiation beamformer $\mathbf{F}_{\mathrm{EM}}$ dynamically adjusts the amplitude responses of each reconfigurable antenna element, facilitating flexible control over energy distribution across diverse spatial directions. Extending the radiation pattern model $\mathbf{f}_{\mathrm{EM}}$ described in Section II-A for a single pattern-reconfigurable antenna, we characterize the collective programmable radiation pattern of the entire antenna array at the BS, comprising $N_{\mathrm{t}}$ reconfigurable antennas. Mathematically, the EM domain radiation beamformer is represented in the following matrix form
@@
where the $m$-th element of $\mathbf{f}_{\mathrm{EM},n}$ represents the radiation gain of the $n$-th antenna in the $m$-th angular direction. To ensure fairness in pattern design, the total energy of each antenna’s radiation pattern is restricted and normalized as $\|\mathbf{f}_{\mathrm{EM},n}\|^{2}=1$. To further delineate the spatial characteristics of reconfigurable antennas, we introduce the virtual angular index channel matrix associated with the $l$-th propagation path for the $k$-th UE
$\mathbf{H}_{\mathrm{EM},k,l}\triangleq\mathbf{I}_{N_{\mathrm{t}}}\otimes\mathbf{h}_{\mathrm{EM},k,l}\in\{0,1\}^{MN_{\mathrm{t}}\times N_{\mathrm{t}}},~\forall k,l,$ (5)
@@
where $\mathbf{h}_{\mathrm{EM},k,l}\in\{0,1\}^{M}$ explicitly specifies the spatial angles associated with the $k$-th UE. Each $\mathbf{h}_{\mathrm{EM},k,l}$ contains exactly one non-zero element, indicating the angular direction of the $l$-th propagation path for the $k$-th UE among $M$ angular sampling points. Consequently, the actual radiation gains of all antennas $\mathbf{G}_{\mathrm{EM},k,l}\in\mathbb{R}^{N_{\mathrm{t}}\times N_{\mathrm{t}}}$ can be formulated as
$\mathbf{G}_{\mathrm{EM},k,l}$ $=\mathbf{H}_{\mathrm{EM},k,l}^{H}\mathbf{F}_{\mathrm{EM}}$ (6)
$\triangleq\mathrm{diag}\{G_{\mathrm{EM},k,l,1},\ldots,G_{\mathrm{EM},k,l,N_{\mathrm{t}}}\},~\forall k,l,$
@@
which is a diagonal matrix with the $n$-th entry $G_{\mathrm{EM},k,l,n}\triangleq\mathbf{h}_{\mathrm{EM},k,l}^{H}\mathbf{f}_{\mathrm{EM},n}$ representing the actual radiation gain of the $n$-th antenna towards the $l$-th path of the $k$-th UE. In summary, the matrices $\mathbf{F}_{\mathrm{EM}}$ and $\mathbf{H}_{\mathrm{EM},k,l}$ mathematically capture the additional DoF afforded by reconfigurable antennas, enabling highly adaptive radiation beamforming in the EM domain.
@@
RF domain: The analog beamformer $\mathbf{F}_{\mathrm{RF}}$ controls the phases of the transmitted signals across antennas, ensuring coherent signal combination and directional beamforming toward desired UEs. We adopt a dynamic-subarray analog beamforming architecture, wherein each RF chain adaptively connects to distinct subsets of antennas through a switching network. This architecture optimally balances system efficiency and communication performance by exploiting dynamic antenna selection. Assuming the number of RF chains $N_{\mathrm{RF}}$ meets or exceeds the number of UEs $K$, i.e., $N_{\mathrm{RF}}\geq K$. The analog beamformer employs a set of low-resolution PSs, each following a constant modulus constraint with quantized phase control of $B$ bits. The discrete phase set of PSs is given by $\mathcal{F}\triangleq\{\frac{1}{\sqrt{N_{\mathrm{t}}}}e^{\jmath\frac{2\pi b}{2^{B}}}|b=0,1,\ldots,2^{B}-1\}$. To facilitate dynamic-subarray beamforming, we define the analog beamformer matrix as $\mathbf{F}_{\mathrm{RF}}\in\{\mathcal{F},0\}^{N_{\mathrm{t}}\times N_{\mathrm{RF}}}$, which implies that if the $l_{\mathrm{RF}}$-th RF chain is connected to the $n$-th antenna, the corresponding phase-shift takes a non-zero value, i.e., $\mathbf{F}_{\mathrm{RF}}(n,l_{\mathrm{RF}})\in\mathcal{F}$; if not, it is set as zero, i.e., $\mathbf{F}_{\mathrm{RF}}(n,l_{\mathrm{RF}})=0$. Moreover, to ensure non-overlapping subarrays, each antenna is dynamically assigned to only one RF chain, meaning that the analog beamformer matrix is restricted to have only one non-zero element per row, i.e., $\|\mathbf{F}_{\mathrm{RF}}(n,:)\|_{0}=1$.